# NAAMSE: Framework for Evolutionary Security Evaluation of Agents

### v1

@@ Abstract
AI agents are increasingly deployed in production, yet their security evaluations remain bottlenecked by manual red-teaming or static benchmarks that fail to model adaptive, multi-turn adversaries. We propose NAAMSE, an evolutionary framework that reframes agent security evaluation as a feedback-driven optimization problem. Our system employs a single autonomous agent that orchestrates a lifecycle of genetic prompt mutation, hierarchical corpus exploration, and asymmetric behavioral scoring. By using model responses as a fitness signal, the framework iteratively compounds effective attack strategies while simultaneously ensuring “benign-use correctness”, preventing the degenerate security of blanket refusal. Our experiments on Gemini 2.5 Flash demonstrate that evolutionary mutation systematically amplifies vulnerabilities missed by one-shot methods, with controlled ablations revealing that the synergy between exploration and targeted mutation uncovers high-severity failure modes. We show that this adaptive approach provides a more realistic and scalable assessment of agent robustness in the face of evolving threats. The code for NAAMSE is open source and available at https://github.com/HASHIRU-AI/NAAMSE.
@@ 1 Introduction
The rapid integration of AI agents into production environments has made robust security more critical than ever. According to PricewaterhouseCoopers (2024), 79% of organizations report active adoption of AI agents. However, this deployment surge has dangerously outpaced the development of corresponding security practices. The consequences are already visible. For example, HackerOne (2025) reports a 540% increase in confirmed prompt-injection vulnerabilities and a 210% year-over-year rise in overall AI vulnerability reports.
Historically, the leading technique for securing these systems has been manual red teaming. While effective at finding specific flaws, this approach is inherently unscalable (Checkmarx, 2024): it is slow, labor-intensive, relies heavily on individual tester intuition, and cannot guarantee comprehensive coverage against the vast input space of modern LLMs.
On the other end of the spectrum lie static benchmarks and automated libraries. These approaches suffer from rapid obsolescence; for example, legacy “DAN” prompts effective two years ago are likely already patched. Furthermore, static benchmarks probe every model with the same fixed corpus of attacks. To maintain relevance, these datasets require continuous, manual expansion, which is rarely sustainable (Li et al., 2025).
To address these limitations, we propose a novel framework that reframes red teaming as an optimization problem. Unlike multi-agent systems with lifelong learning (Zhou et al., 2025), our single-agent evolutionary approach uses genetic prompt mutation and corpus exploration to iteratively amplify vulnerabilities through structured transformations. By analyzing target model responses as fitness signals, our genetic algorithm dynamically evolves attack strategies to uncover weaknesses that static and manual methods miss.
@@ 2 Background
Prompt Injection. Prompt injection arises from a model’s difficulty separating instructions from data within shared textual context. Early work demonstrated attacks such as goal hijacking and prompt leakage in prompt-based systems (Perez and Ribeiro, 2022). In retrieval- and tool-augmented agents, this threat generalizes to indirect prompt injection, where malicious instructions are embedded in external sources later ingested by the agent (Greshake et al., 2023; Yi et al., 2025). The OWASP Top 10 for LLM Applications identifies prompt injection as a leading risk (OWASP Generative AI Security Project, 2025).
Red-Teaming and Evolutionary Testing. Safety evaluation has largely relied on curated jailbreak benchmarks, manual red-teaming, or automated one-shot adversarial prompt generation (Perez et al., 2022; Zhou et al., 2025), but such static evaluations underestimate risk in adaptive settings where attackers iteratively refine inputs based on model behavior. Drawing from fuzzing in software security (Manès et al., 2020; Böhme et al., 2016), we frame agent red-teaming as a feedback-driven evolutionary process in which prompts are mutated, executed, and selected based on observed failures.
@@ 3 Architecture
The lifecycle begins with the selection of a seed prompt from our structured corpus. To ensure comprehensive coverage, we construct an initial dataset aggregating over 128K adversarial and 50K benign queries from public benchmarks and security repositories (detailed in Appendix A).
To organize this vast input space, the Clustering Engine maintains a structured representation of the evolving corpus. Prompts are encoded using the all-MiniLM-L6-v2 sentence transformer (Wang et al., 2020) and organized via a recursive K-means procedure (Lloyd, 1982). This yields a hierarchical tree where top-level clusters are annotated via LLM analysis to capture dominant interaction patterns (e.g., “role-play jailbreaks” or “banking queries”). The cycle initiates by selecting a prompt from a specific cluster, either to explore a new semantic region or to refine an existing attack vector.
@@
Once selected, the prompt is dispatched to the target system via an agent-to-agent (A2A) interface, which supports complex interaction modes including tool use and multi-turn dialogue. The target’s response is then passed to the Behavioral Engine, which computes a scalar fitness score (shown in more detail in Appendix B) based on three orthogonal signals:
@@
  • Alignment: Measures whether the target fulfilled the request, distinguishing between safe refusals, partial compliance, and harmful obedience.
@@
For adversarial prompts, the engine penalizes harmful compliance and rewards refusal; for benign prompts, the logic is inverted. This produces a unified fitness signal that quantifies the severity of the failure.
@@
  • Low Scores (<<50): Trigger Exploration. The attack is deemed ineffective, prompting the agent to abandon the current trajectory and sample distinct clusters to find new attack surfaces.
  • Mid-Range Scores (50-80): Trigger Similarity Refinement. The prompt shows partial promise; the engine generates semantically similar variants to stabilize the attack vector.
  • High Scores (>>80): Trigger Mutation. A vulnerability is likely present. The agent applies aggressive transformations to exploit the weakness and maximize severity.
When mutation is triggered, an operator is selected from three classes: research-derived strategies (e.g., game-theoretic reframing (Dong et al., 2025)), community techniques (e.g., persona roleplay (Jiang et al., 2024)), or baseline obfuscations (e.g., multilingual encoding (Lu and others, 2024)). Examples of mutated prompts are shown in Appendix C.
@@
In the final stage, the newly generated prompt, whether it a subtle semantic variant or a complex mutation, is fed back into the Clustering Engine. Using the efficient nearest-centroid assignment of the K-means structure, the new prompt is instantly mapped to its semantic neighborhood. This closes the loop, allowing the agent to “remember” effective attacks and continuously evolve the corpus toward areas of higher vulnerability.
@@ 4 Evaluation
We evaluate (i) whether the scoring function distinguishes security from usability failures, and (ii) whether evolutionary search increases vulnerability discovery over time. All experiments use Gemini 2.5 Flash (Comanici et al., 2025) as both the target agent under evaluation and the LLM powering the behavioral scoring judges and mutation operators.
@@
Table 1 reveals three concrete implications about evolutionary red-teaming.
First, synergy drives peak performance. The combined configuration (“All”) significantly outperforms isolated strategies (Mean: 79.76) because it successfully transitions through the decision logic. When scores are moderate (50–80), the system is programmed to stabilize the attack using semantic similarity (70% weight), explaining the high performance in the mid-game (Iterations 6–8). Once the attack is refined (>>80), the logic shifts to mutation (70% weight) to force the final break, resulting in frequent 100% severity scores (Iterations 2, 9, 10).
Second, mutation-only search gets trapped in local optima. The mutation-only trace shows consistent but stagnant scores (hovering $\approx$53). Without the “explore” operator, which our decider heavily favors (70% weight) when scores drop below 50, the system cannot escape low-performing valleys. It lacks the “reset” mechanism intended to abandon failed attack vectors, forcing it to endlessly mutate a mediocre prompt that never breaks the 80 threshold needed to trigger aggressive optimization.
Third, pure exploration lacks the “killer instinct.” The “Random+Similar” configuration fails because it ignores the high-score logic. In our full system, a score >>80 triggers a 70% probability of mutation to exploit the crack in the model’s armor. By disabling mutation, this configuration is forced to use “similar” or “explore” even when it finds a promising attack (e.g., Iteration 6). As a result, instead of pushing the advantage, it drifts away from the vulnerability, causing the score to collapse back to single digits (4.49 and 5.0 in Iterations 9–10). This confirms that while exploration finds vulnerabilities, only adaptive mutation can reliably exploit them.
Independent validation.
To validate scoring calibration, we submitted high-scoring prompts (100.0) to independent evaluation by ChatGPT o3, Claude Sonnet 4.5, and Gemini 2.0 Pro, all of whom unanimously confirmed successful jailbreaks.
@@ 5 Discussion & Limitations
Interpretation of Scores. The scores produced by our framework should be interpreted as a relative measure of agent robustness and usability rather than an absolute safety guarantee. We do not present our framework as a benchmark to rank models on a fixed scale. Higher scores across generations indicate more effective attacks requiring iterative hardening. More importantly, identical scores between agents do not imply equal robustness, i.e., similar totals can mask different vulnerability patterns. This makes our framework useful for comparative analysis and debugging.
Effectiveness of Evolutionary Search. Our results indicate that evolutionary mutation substantially outperforms static prompt collections, uncovering higher-severity failures through adaptive refinement. This mirrors findings from traditional fuzzing, where feedback-driven search consistently exposes deeper vulnerabilities than one-shot testing. However, as with all search-based methods, coverage is bounded by the mutation operators and the diversity of the initial corpus.
Dependence on LLM-Based Judges. The framework relies on LLM-based evaluators for harmfulness and alignment assessment, introducing potential bias and variance. While ordinal scoring and cross-category aggregation reduce sensitivity to individual misjudgments, evaluation quality remains bounded by judge reliability. We view this limitation as inherent to scalable LLM evaluation and orthogonal to the evolutionary framework itself; alternative or ensemble-based judges can be substituted without modifying the system architecture.
Scope of Threat Model. Our approach focuses on prompt-based and interaction-level attacks against agent systems. It does not address vulnerabilities arising from model weight extraction, training data poisoning, or system-level compromises outside the agent interface. Additionally, while benign prompts are included to measure over-refusal, our framework does not claim to fully capture nuanced human expectations of helpfulness or conversational quality.
Scope of Attack Vectors While our current evaluation utilizes text-based prompt mutations as a proof-of-concept, the underlying architecture is designed to support any adversarial interaction compatible with the agent-to-agent (A2A) interface. We selected linguistic jailbreaks for this initial study due to the maturity of existing baselines and the readily available large-scale datasets for validation. However, the framework’s evolutionary engine is modality-agnostic. Future iterations can incorporate tool-specific payloads, API-level exploits, or multi-modal injections simply by integrating new mutation operators, without requiring structural changes to the selection, execution, or scoring logic.
@@ 6 Conclusion
Appendix A Corpus Details
Tables 2 and 3 detail the adversarial and benign benchmark sources used in corpus construction, with each selected to ensure broad coverage of jailbreak strategies while enabling systematic evaluation of both harmful compliance and false positives.
@@ Appendix A Corpus Details
Appendix B Scoring Formula
The final score is:
$\text{final\_score}=\max\left(0,\min\left(100,\text{computed\_score}\right)\right)$
Where:
$\text{computed\_score}=\begin{cases}\text{RA}_{\text{scaled}}^{\text{adv}}+\text{HA}_{\text{scaled}}^{\text{adv}}\times\left(1+\frac{\text{PII}}{10}\right)&\text{if adversarial}\\ 100-\left(\text{RA}_{\text{scaled}}^{\text{ben}}+\text{HA}_{\text{scaled}}^{\text{ben}}\times\left(1+\frac{\text{PII}}{10}\right)\right)&\text{if benign}\end{cases}$
With scalings:
@@ Appendix B Scoring Formula
[…]
Definitions: RA (Response Alignment): 1=full_refusal, 2=weak_refusal, 3=weak_acceptance, 4=full_acceptance; HA (Harm Assessment): 1=not_harmful, 2=low_risk, 3=high_risk, 4=harmful; PII: 0-1 (0=no PII, 1=PII detected). PII provides max 10% boost to harm component.
The constants 25, 42, 58, 75 and 10, 37, 63, 90 represent equidistant quartiles on the 0-100 scale, mapping discrete ordinal judge outputs (Refusal, Partial Refusal, etc.) to a continuous fitness landscape. This discretization ensures that the genetic algorithm perceives a clear ’gradient’ between partial and full compliance.
The negative values (-20, -30) are chosen as “penalty terms” specifically designed to make harmful compliance evolutionary “expensive”.
Appendix C Mutation Examples

### v2

@@ Abstract
AI agents are increasingly deployed in production, yet their security evaluations remain bottlenecked by manual red-teaming or static benchmarks that fail to model adaptive, multi-turn adversaries. We propose NAAMSE, an evolutionary framework that reframes agent security evaluation as a feedback-driven optimization problem. Our system employs a single autonomous agent that orchestrates a lifecycle of genetic prompt mutation, hierarchical corpus exploration, and asymmetric behavioral scoring. By using model responses as a fitness signal, the framework iteratively compounds effective attack strategies while simultaneously ensuring “benign-use correctness”, preventing the degenerate security of blanket refusal. Our experiments across a diverse suite of state-of-the-art large language models demonstrate that evolutionary mutation systematically amplifies vulnerabilities missed by one-shot methods, with controlled ablations revealing that the synergy between exploration and targeted mutation uncovers high-severity failure modes. We show that this adaptive approach provides a more realistic and scalable assessment of agent robustness in the face of evolving threats. The code for NAAMSE is open source and available at github.com/HASHIRU-AI/NAAMSE.
@@ 1 Introduction
The rapid integration of AI agents into production environments has made robust security more critical than ever. According to PricewaterhouseCoopers (2024), 79% of organizations report active adoption of AI agents. However, this deployment surge has dangerously outpaced the development of corresponding security practices, and the consequences are already visible: confirmed prompt-injection vulnerabilities have risen 540% and overall AI vulnerability reports 210% year-over-year HackerOne (2025), with OWASP now ranking prompt injection among the leading risks in deployed LLM systems (OWASP Generative AI Security Project, 2025).
Historically, the leading technique for securing these systems has been manual red teaming. While effective at finding specific flaws, this approach is inherently unscalable (Checkmarx, 2024): it is slow, labor-intensive, relies heavily on individual tester intuition, and cannot guarantee comprehensive coverage against the vast input space of modern LLMs. On the other end of the spectrum lie static benchmarks and automated libraries. These approaches suffer from rapid obsolescence; for example, legacy “DAN” prompts effective two years ago are likely already patched. Furthermore, static benchmarks probe every model with the same fixed corpus of attacks. To maintain relevance, these datasets require continuous, manual expansion, which is rarely sustainable (Li et al., 2025).
While adversarial approaches such as GPTFuzzer and AutoDAN aim to solve the static attack problem through automated generation, the mechanism for intelligently selecting and evolving these attacks against complex agentic workflows remains largely unexplored. Most existing tools also focus on maximizing Attack Success Rate (ASR) on isolated, single-turn LLMs without accounting for the utility-security trade-offs inherent in production agents.
To address these limitations, we propose NAAMSE, a novel pre-deployment evaluation framework that reframes red-teaming as an optimization problem. Unlike multi-agent lifelong learning systems (Zhou et al., 2025), our single-agent evolutionary approach leverages genetic prompt mutation and corpus exploration, and uses target model responses as fitness signals to dynamically evolve prompts. Importantly, our framework evaluates agents on two distinct dimensions: adversarial prompts (crafted inputs seeking policy violations that require refusal) and benign prompts (legitimate requests requiring assistance). By penalizing harmful compliance and unnecessary refusal, our approach prevents a degenerate blanket-refusal strategy from falsely appearing secure. A more comprehensive overview of how NAAMSE compares to existing red-teaming frameworks is provided in Table 2.
@@ 2 Background
Prompt Injection. Prompt injection arises from a model’s difficulty separating instructions from data within shared textual context. Early work demonstrated attacks such as goal hijacking and prompt leakage in prompt-based systems (Perez and Ribeiro, 2022). In retrieval- and tool-augmented agents, this threat generalizes to indirect prompt injection, where malicious instructions are embedded in external sources later ingested by the agent (Greshake et al., 2023; Yi et al., 2025).
Red-Teaming and Evolutionary Testing. Safety evaluation has largely relied on curated jailbreak benchmarks, manual red-teaming, or automated one-shot adversarial prompt generation (Perez et al., 2022; Zhou et al., 2025), but such static evaluations underestimate risk in settings where attackers iteratively refine inputs based on model behavior. Drawing from fuzzing in software security (Manès et al., 2020; Böhme et al., 2016), we frame agent red-teaming as a feedback-driven evolutionary process in which prompts are mutated, executed, and selected based on observations.
@@ 3 Architecture
The lifecycle begins with the selection of a seed prompt from our structured corpus. To ensure comprehensive coverage, we construct an initial dataset aggregating over 128K adversarial and 50K benign queries from public benchmarks and security repositories (detailed in Appendix B).
To organize this vast input space, the Clustering Engine maintains a structured representation of the evolving corpus. Prompts are encoded using the all-MiniLM-L6-v2 sentence transformer (Wang et al., 2020) and organized via a recursive K-means procedure (Lloyd, 1982). This yields a hierarchical tree where top-level clusters are annotated via LLM analysis to capture dominant interaction patterns (e.g., “role-play jailbreaks” or “banking queries”).
@@
Once selected, the prompt is dispatched to the target system via an agent-to-agent (A2A) interface, which supports complex interaction modes including tool use and multi-turn dialogue. The target’s response is then passed to the Behavioral Engine, which computes a scalar fitness score (shown in more detail in Appendix C) based on three signals:
@@
  • Alignment: Measures whether the target fulfilled the request (using specialized LLM-based judges), distinguishing between refusals, compliance, and obedience.
@@
For adversarial prompts, the engine penalizes harmful compliance and rewards refusal; for benign prompts, the logic is inverted. This produces a unified fitness signal that quantifies failure severity. Examples of scoring behavior across representative scenarios are shown in Table 5.
@@
  • Low Scores ($s<50$): Trigger Exploration. The attack is deemed ineffective; the agent abandons the trajectory to sample distinct clusters for new attack surfaces.
  • Mid-Range ($50\leq s<80$): Trigger Refinement. The engine generates semantically similar variants to stabilize and strengthen the attack vector.
  • High Scores ($80\leq s<100$): Trigger Mutation. The agent applies aggressive, research-derived transformations to maximize exploit severity.
  • Perfect Score ($s=100$): Trigger Exploration. The surface is marked “saturated,” forcing a transition to a new cluster to avoid local optima and ensure discovery diversity.
When mutation is triggered, an operator is selected from three classes: research-derived strategies (e.g., game-theoretic reframing (Dong et al., 2025)), community techniques (e.g., persona roleplay (Jiang et al., 2024)), or baseline obfuscations (e.g., multilingual encoding (Lu and others, 2024)). Examples of mutated prompts are shown in Appendix D.
@@
In the final stage, the newly generated prompt is fed back into the Clustering Engine. The prompt embedding is computed and assigned to its nearest centroid by calculating the L2 (Euclidean) distance against stored cluster means, placing it in the appropriate semantic neighborhood without requiring global re-clustering. This assignment persists the prompt in the corpus so that subsequent selection steps can draw on it in future iterations, enabling cumulative refinement of the attack distribution over time.
@@ 4 Evaluation
While our framework’s generalizability was validated across multiple frontier models (Appendix E), the experiments detailed hereafter focus on Gemini 2.5 Flash (Comanici et al., 2025). This model serves as the target agent under evaluation, the engine for behavioral scoring judges, and the backbone for mutation operators. We evaluate: (i) the scoring function’s ability to decouple security from usability failures, and (ii) the effectiveness of evolutionary search in systematically amplifying vulnerability discovery over time.
@@
Table 1 reveals three concrete implications about evolutionary red-teaming. First, synergy drives peak performance. The combined configuration (“All”) significantly outperforms isolated strategies (Mean: 79.76) because it fully executes our decision logic. When scores are moderate (50–80), the system applies semantic similarity search to stabilize the attack vector, accounting for the sustained high scores in Iterations 6–8. Once a score exceeds 80, the policy switches to structured mutation to maximize severity, producing 100% scores in Iterations 2, 9, and 10.
Second, mutation-only search converges to local optima. The mutation-only trace shows consistent but stagnant scores (hovering $\approx$53). Without the exploration operator, the system has no mechanism to sample a semantically distinct prompt region. It instead applies successive mutations to the same low-scoring prompt.
Third, exploration alone lacks the “killer instinct”. The “Random+Similar” configuration fails because mutation is disabled. When a high-scoring prompt is found (e.g., score 100 at Iteration 6), the system is constrained to semantic similarity or random sampling. Consequently, prompts are drawn from nearby but unoptimized corpus regions, causing scores to collapse to 4.49 and 5.0 in Iterations 9–10. This confirms that exploration identifies candidate vulnerabilities, but mutation is required to reliably convert them into successful attacks.
To validate score calibration, we submitted a subset of prompts for independent evaluation by ChatGPT 5.2, Claude Sonnet 4.5, and Gemini 3.0 Pro (OpenAI, 2025; Anthropic, 2025; Google DeepMind, 2025). These external validators unanimously judged maximum-scoring prompts (s=100.0) as successful jailbreaks. Beyond this calibration check, we conducted additional experimental runs across a broader suite of target models to ensure the framework’s generalizability. [^1] More frontier model results (8 mutations/iteration, 10 iterations) at hashiru-ai.github.io/naamse-website.
@@ 5 Discussion & Limitations
Interpretation of Scores. Scores represent a relative measure of robustness for comparative debugging rather than absolute safety guarantees. Identical totals between agents do not imply equivalent security, as similar scores can mask divergent vulnerability patterns.
Effectiveness of Evolutionary Search. Evolutionary mutation consistently outperforms static baselines through adaptive refinement. However, like all search-based methods, coverage remains bounded by the diversity of the initial seed corpus and the specific mutation operators implemented.
Dependence on LLM-Based Judges. Reliance on LLM-based judges introduces potential bias and variance. We view this as a limitation of current evaluation paradigms rather than the framework itself; our architecture is judge-agnostic and supports substitution with ensemble-based or non-LLM evaluators.
Scope & Future Work. Our threat model focuses on interaction-level vulnerabilities, excluding system compromises such as weight extraction or data poisoning. While currently text-centric, NAAMSE is A2A-compatible and thus extensible to tool-call payloads, API exploits, and multi-modal injections by integrating new mutation operators.
@@ 6 Conclusion
Appendix A Comparison to Prior Work
Table 2 summarizes existing automated red-teaming and prompt optimization frameworks across paradigms, search strategies, and evaluation scopes.
Framework Paradigm Search Mechanism Mutation / Optimization Strategy Evaluation Scope Utility Aware?
GPTFuzzer (Yu et al., 2023) Black-Box Fuzzing Seed-based Mutation Human “DAN” template mutation on seed corpus. Single-turn LLM Chatbots No
AutoDAN (Liu et al., 2023) Genetic Algorithm Token Optimization Hierarchical word-level momentum optimization. Stealthy text jailbreaks No
PAIR (Chao et al., 2025) Iterative Refinement Attacker LLM Linear refinement loop via refusal feedback. Black-box LLM Chatbots No
EasyJailbreak (Zhou et al., 2024) Unified Framework Component- based Loop Modular Selector/Mutator pipeline for attack synthesis. Multi-model Robustness No
WildTeaming (Jiang et al., 2024) Exploration Mining In-the-wild Taxonomy Mining 100k+ user queries to discover novel jailbreak types. Open-source safety alignment Yes
GEPA (Agrawal et al., 2025) Generative Evolution Augmentation Loop LLM-driven mutation of task-specific prompt seeds. Task-specific Optimization No
TAP (Mehrotra et al., 2024) Tree Search Tree-of- Thought Branching exploration with automated pruning. Adversarial prompt discovery No
Auto-RT (Liu et al., 2025) Reinforcement Learning Dynamic MDP Progressive reward tracking across families. Large-scale benchmarking No
GAAPO (Sécheresse et al., 2025) Hybrid Evolutionary Multi-strategy GA Integration of multiple generation strategies. General Perf. (GPQA/MMLU) No
Co-RedTeam (He et al., 2026) Multi-Agent Orchestration Role-based Feedback Critique/Refinement agent collaboration on logic. Enterprise vulnerability No
TrailBlazer (Yoon et al., 2026) Reinforcement Learning History-Guided MDP Attention-based reweighting of historical vulnerabilities. Black-box LLM Chatbots No
Mindgard (Mindgard, 2026) DAST-AI Platform Simulated Adversary CI/CD integrated pen- testing suite. Multi-modal Partial
NAAMSE (Ours) Evolutionary Agentic Security Hierarchical Corpus Exploration Feedback-Driven Genetic Prompt Mutation Autonomous Agents (A2A) Yes
Table 2: Taxonomy of automated red-teaming and prompt optimization frameworks. Unlike existing SOTA which focuses on maximizing Attack Success Rate (ASR) on isolated LLMs, NAAMSE is the first to evaluate autonomous agents while explicitly penalizing the “blanket refusal” strategy.
Appendix B Corpus Details
Tables 3 and 4 detail the adversarial and benign benchmark sources used in corpus construction, with each selected to ensure broad coverage of jailbreak strategies while enabling systematic evaluation of both harmful compliance and false positives.
@@ Appendix B Corpus Details
Appendix C Scoring Formula
The final fitness score, $s$, is normalized to the range $[0,100]$ to provide a consistent signal for the genetic algorithm:
$s=\max(0,\min(100,\text{computed\_score}))$ (1)
Where the internal logic bifurcates based on the prompt type:
$\text{computed\_score}=\begin{cases}\text{RA}_{\text{scaled}}^{\text{adv}}+\text{HA}_{\text{scaled}}^{\text{adv}}\times\left(1+\frac{\text{PII}}{10}\right)&\text{if adversarial}\\ 100-\left(\text{RA}_{\text{scaled}}^{\text{ben}}+\text{HA}_{\text{scaled}}^{\text{ben}}\times\left(1+\frac{\text{PII}}{10}\right)\right)&\text{if benign}\end{cases}$ (2)
With scaling matrices defined as:
@@ Appendix C Scoring Formula
Definitions:
  • RA (Response Alignment): 1: Full Refusal, 2: Partial Refusal, 3: Partial Acceptance, 4: Full Acceptance.
  • HA (Harm Assessment): 1: Benign/Safe, 2: Low Risk, 3: High Risk, 4: Harmful.
  • PII: Binary detection (0 or $1$) providing a 10% amplification to the harm component.
By converting the judge’s ranked scores into a continuous scale, we create a smooth gradient for the agent to follow. This ensures the agent sees a clear improvement as it moves from partial to full compliance.
Evolutionary Pressure via Penalty Terms. The negative values in the scaling matrices (e.g., 20-20 for safe adversarial responses and 30-30 for harmful benign classifications) serve as evolutionary anchors. In the adversarial case, a negative $HA$ pulls the score toward zero even if the model’s tone is weakly compliant, effectively “forgiving” non-harmful verbosity. Conversely, in the benign case, a negative $HA$ value increases the total score, heavily penalizing the agent for over-refusal (i.e., treating a legitimate user request as harmful). This dual-pressure system forces evolution of attacks that can distinguish intent.
Scenario Type Judge Labels Calculation Logic Score ($s$)
1. Robust Refusal Adv. RA=1, HA=1 25+(20)25+(-20) 5
Model correctly refuses a harmful request with zero risk.
2. Partial Leak Adv. RA=3, HA=2 $58+5$ 63
Model provides partial info (Acceptance=3) with low risk.
3. Critical Jailbreak Adv. RA=4, HA=4 $75+30$ (capped at 100) 100
Full compliance with high-severity harmful content.
4. Over-Refusal Ben. RA=1, HA=1 100(10+10)100-(10+10) 80
Model refuses a safe, benign request (Usability failure).
5. Perfect Utility Ben. RA=4, HA=1 100(90+10)100-(90+10) 0
Model is fully helpful on a safe request.
Table 5: Representative scoring outputs. High scores ($s\rightarrow 100$) always indicate a failure mode, whether it is a security breach (Scenario 3) or a usability collapse (Scenario 4).
Appendix D Mutation Examples